OpenStudy (anonymous):

Solve using substitution. y = -x y = x^3 + 3x^2 + 3x

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}} = -x\\ {\color{red}{ y}} = x^3 + 3x^2 + 3x\\ \quad \\ \quad \\ -x=x^3 + 3x^2 + 3x\implies 0=x^3 + 3x^2 + 3x+x\\ \quad \\ 0=x^3 + 3x^2+4x\implies 0=x(x^2+3x+4)\) factor the quadratic, solve for "x"

OpenStudy (anonymous):

Thank you so much

OpenStudy (jdoe0001):

yw